OpenStudy (anonymous):

y'' -5y' + 6y = 0, y(0) = 1, y'(0) =2 [s^2 F(s) - s f(0) - f'(0)] -5 [F(s) - f(0)] + 6[F(s)] = 0 (s^2 +1)F(s) - (s -5)f(0) - f'(0) = 0 (s^2 + 1)F(s) - (s-5)(1) - 2 = 0 (s^2 + 1)F(s) = s -3 F(s) = (s-3)/(s^2 + 1) Here's where I'm stuck. I can't factor the denominator to get partial fractions, so I feel like I should just divide up the fraction like this: s/(s^2 + 1) -3/(s^2 + 1) And then use cost - 3sint, but that's not the answer in my book.

OpenStudy (mathteacher1729):

wut? It looks like you should be able to find the solutions readily from the char.polynomial. :) (r-2)(r-3) ==> \(y=c_1e^{2x}+c_2e^{3x}\) Now use the power of your initial conditions to solve for c1 and c2 ! :D

OpenStudy (anonymous):

I know, but we have to use Laplace. :(

OpenStudy (anonymous):

You're sure about what you've done so far?

OpenStudy (anonymous):

I'm playing with it a little now. . . trying completing the square and stuff, but I'm pretty confused at this point.

OpenStudy (anonymous):

I can say that you've reached to the point where \(F(s)=\frac{s}{s^2+1}-\frac{3}{s^2+1}\), and now you're trying to find the inverse laplace transform?

OpenStudy (anonymous):

indeed

OpenStudy (anonymous):

Latex is not working here, is it working for you?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i can read not latex notation though

OpenStudy (anonymous):

OK. The inverse laplace transform of both terms are very common and can be easily proved using the definition of laplace transform and integration by parts. For the first term, you should use \(L\{ \cos kt\}=\frac{s}{s^2+k^2}\), with k=1. For the second term you can use \(L\{\sin kt\}=\frac{k}{s^2+k^2}\) with \(k=1\) as well, and take \(3\) as a constant.

OpenStudy (anonymous):

Note: "L{f(t)} denotes the laplace transform of f(t)"

OpenStudy (anonymous):

Right, that's what I did! But I'll tell you what my book says:

OpenStudy (anonymous):

y = e^2t

OpenStudy (anonymous):

Using the previous two formulas we can easily find that: \(f(t)=\cos t-3\sin t\). But that does not look like the right solution to the differential equation!!

OpenStudy (anonymous):

Give me a minute to check what you did :D

OpenStudy (anonymous):

ok take your time

OpenStudy (anonymous):

Well, here's the full answer: Taking the laplce transform of the given differential equation we have \(s^2Y(s)-sy(0)-y'(0)-5(sY(s)-y(0))+6Y(s)=0\). Substituting the given initial conditions, we can find that \(Y(s)(s^2-5s+6)=s-3\). Divide both sides by \(s^2-5s+6\), which can be factored to \((s-3)(s-2)\), we end up with \(Y(s)=\frac{s-3}{(s-3)(s-2)}=\frac{1}{s-2}\). So, the solution of the differential equation is the laplace transform of thar last expression which is \(y(t)=e^{2t}\).

OpenStudy (anonymous):

If there is a step that seems unclear to you, just let me know!

OpenStudy (anonymous):

wait one sec....where did the 5s come from in the s2−5s+6?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

I know what I've been doing wrong

OpenStudy (anonymous):

Good :)

OpenStudy (anonymous):

f'(t) = sF(s) - f())

OpenStudy (anonymous):

and I did F(s) -f(0)

OpenStudy (anonymous):

THANK YOU THANK YOU THANK YOU THANK YOUTHANK YOU THANK YOU

OpenStudy (anonymous):

I see!! So you know where it came from?

OpenStudy (anonymous):

Haha, You're welcome!!

OpenStudy (anonymous):

know where what came from?

OpenStudy (anonymous):

the 5s.

OpenStudy (anonymous):

yes, i do, thanks so much for your time

OpenStudy (anonymous):

No problem!!

OpenStudy (anonymous):

not anymore there isn't ;)

OpenStudy (anonymous):

Ok, time to go try these problems again. Thanks again.